OpenStudy (aaronandyson):

Find p if the quadratic equation 4x^2 + (p-2)x + 1 = 0 has equal rooots.

OpenStudy (aaronandyson):

@GirlgoyleH. @ganeshie8 @Gabylovesyou @ilovewolf

OpenStudy (aaronandyson):

@SolomonZelman

OpenStudy (paki):

@Nnesha help kar lo ess ki...

OpenStudy (girlgoyleh.):

im sorry im horrid at math

OpenStudy (aum):

A quadratic equation will have equal roots when the discriminant is zero.

OpenStudy (aum):

4x^2 + (p-2)x + 1 = 0 Discriminant = b^2 - 4ac = 0 or b^2 = 4ac That is, (p-2)^2 = 4*4*1 = 16 (p-2)^2 - 16 = 0 (p-2)^2 - 4^2 = 0 (p-2+4)(p-2-4) = 0 (p+2)(p-6) = 0 p = -2, +6

OpenStudy (jhannybean):

Vieta's Formulas?! @ganeshie8

OpenStudy (aaronandyson):

Please someone help I'm totally confused. :(

OpenStudy (raden_zaikaria):

im got two answer p=7 or p=6

OpenStudy (fibonaccichick666):

ok, question, if instead of p-2 you just had an x there, could you find x?

OpenStudy (fibonaccichick666):

there would be two values you should find for x

OpenStudy (fibonaccichick666):

@AaronAndyson

OpenStudy (jhannybean):

Notice that in @aum's problem, he simplify reinstated the quadratic formula by comparison: \[\color{red}ax^2+\color{blue}bx+\color{orange}c=0 \]\[\color{red}4x^2 +\color{blue}{(p-2)}x+\color{orange}1=0\]The discriminant of the quadratic formula tells you not only how many roots there are, but also helps you find out what they are, so now just compare your equation to the quadratic, plug it into the formula for the discriminant and solve to find your roots :)

OpenStudy (jhannybean):

The discriminant is stated as: \(b^2-4ac\)

OpenStudy (aaronandyson):

I did not get what you said @FibonacciChick666

OpenStudy (aaronandyson):

What ? @Jhannybean

OpenStudy (fibonaccichick666):

ok, if you have \[4x^2 + rx + 1 = 0\] can you find r? Sorry, bout the x, I didn't realize that was your variable

OpenStudy (aaronandyson):

? What ?

OpenStudy (jhannybean):

You're not even trying, it seems.

OpenStudy (fibonaccichick666):

can you factor \[4x2+rx+1=0\] please

OpenStudy (fibonaccichick666):

sorry, \[4x^2+rx+1=0\]

OpenStudy (aaronandyson):

@Jhannybean the thing I'm doing this after a very long period of time so it seems like I have forgotten every thing.

OpenStudy (jhannybean):

So you're unfamiliar with factoring complex quadratics?

OpenStudy (aaronandyson):

It can be said like that.

OpenStudy (aum):

4x^2 + (p-2)x + 1 = 0 divide by 4: x^2 + (p-2)/4*x + 1/4 = 0 --- (1) Assume the equal roots are r and r Then (x-r)(x-r) = 0 x^2 - 2rx + r^2 = 0 ----- (2) From (1) and (2): x^2 + (p-2)/4*x + 1/4 = x^2 - 2rx + r^2 Equate the corresponding coefficinets: (p-2)/4 = 2r ---- (3) r^2 = 1/4 implies 4r^2 = 1 ---- (4) Square (3): (p-2)^2 / 16 = 4r^2 From (4): (p-2)^2 / 16 = 1 (p-2)^2 = 16 Take square root: p-2 = +4 or p-2 = -4 p = 6 or p = -2

OpenStudy (aum):

There is a negative sign missing above but it should not affect the final answer (p-2)/4 = -2r ---- (3) (-2r and nor 2r)

OpenStudy (aaronandyson):

pfff nvm

OpenStudy (raden_zaikaria):



OpenStudy (raden_zaikaria):